OpenStudy (anonymous):

The product of two consecutive positive even numbers is 960. What are the numbers? (Enter solutions from smallest to largest.)

OpenStudy (cwrw238):

n (n + 2) = 960 n^2 + 2n - 960 = 0 you need to solve this for n can you do thid?

OpenStudy (cwrw238):

the 2 numbers are 30 and 32

OpenStudy (cwrw238):

ok?

OpenStudy (cwrw238):

30 * 32 = 30 * 30 + 30 * 2 = 900 + 60

OpenStudy (anonymous):

Thank you

OpenStudy (cwrw238):

yw